OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundredth. 5x + 14 = 26

OpenStudy (anonymous):

\[5x+14=26\] subtract 14 from both sides \[5x+14-14=26-14\] \[5x=12\] divide both sides by 5 \[5x/5=12/5\] The solution is 12/5 which decimal form is 2.4